OpenStudy (anonymous):

Please show work so I can understant it. I have my first test tomorrow morning. 4x+3=-3+2x+14

OpenStudy (watchmath):

Simplify the -3+14 on the right 4x+3=2x+11 Add -3 to both sides 4x=2x+8 Add -2x to both sides 2x=8 Divide by 2 x=4

OpenStudy (anonymous):

this is simple work really. purple math.com/

OpenStudy (anonymous):

try mathway.com for showing all the steps.